the length of a rectangular garden is 7 feet longer tha its width. the perimeter of the garden is 38 feet find the width and length of the garden?

Answer by checkley77(12569)

L=W+9
2L+2W=38
2(W+9)+2W=38
2W+18+2W=38
4W=38-18
4W=20
W=20/4
W=5 ANS. FOR THE WIDTH.
L=5+9
L=14 ANS. FOR THE LENGTH.
PROOF:
2*14+2*5=38
28+10=38
38=38